OpenStudy (sonyalee77):

Solve the inequality and then graph its solution: 15<=3/4(f+4)

OpenStudy (sonyalee77):

@texaschic101 @wolfe8 @eliassaab

OpenStudy (sonyalee77):

@Jack1 @Bacon892003

OpenStudy (sonyalee77):

@RANE

OpenStudy (sonyalee77):

@bakonloverk

OpenStudy (sonyalee77):

Already done with it....

OpenStudy (anonymous):

15<=3/4(f+4) 60 <= 3(f+4) 60 <= 3 f +12 60-12 <= 3 f 48 <= 3 f f >= 16